OpenStudy (anonymous):

How do I add sqrt(3x+7) to sqrt(3x-7)?

OpenStudy (anonymous):

\[\sqrt{3x+7}+\sqrt{3x+7}\] x + x =?

OpenStudy (anonymous):

@keeponbleeding you there

OpenStudy (anonymous):

The second one is 3x-7 not +7 but x+x =2x...

OpenStudy (anonymous):

\[\sqrt{3x+7}+\sqrt{3x-7}\] you cant add them unless you have a value for x

OpenStudy (anonymous):

well actually the question was a function: f(x)=sqrt(3x+7) g(x)= sqrt(3x-7)? Find (f + g)(x). So I thought Id add them…

OpenStudy (anonymous):

correct just put the "+" between them and you are good to go

OpenStudy (anonymous):

ohhh. k, its that easy...thanks! :)